OpenStudy (anonymous):

Can someone double check: 1.lim x-->0 : x^2+2x/x^4 cancel x^2 and x^4 = 2x/x^2 direct substitute, lim DNE or 0 2. Direct substitution: lim x-->pi/2 : (2e^xsinx) = pi/2

OpenStudy (anonymous):

@SithsAndGiggles ?

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{x^2+2x}{x^4}=\lim_{x\to0}\frac{x+2}{x^3}\] The limit doesn't exist. For the second limit, directly substituting gives you \[\lim_{x\to\pi/2}2e^x\sin x=2e^{\pi/2}\sin \frac{\pi}{2}=2e^{\pi/2}\]

OpenStudy (anonymous):

for 2. I put in 2e^pi/2 but it was the incorrect answer. my other choices are 0,1, or pi/2 @SithsAndGiggles

OpenStudy (anonymous):

Not sure what to tell you: http://www.wolframalpha.com/input/?i=Limit%5B2*Exp%5Bx%5D*Sin%5Bx%5D%2Cx-%3Epi%2F2%5D Maybe you mean a different function?

OpenStudy (anonymous):

I'll choose it again, thanks

OpenStudy (anonymous):

It's correct, my bad